OpenStudy (anonymous):

1+1= (\_/) (O_O)o0O[ ..H E L P .. ] .(v v) >_ >'

hartnn (hartnn):

asked thousand times here.... its 10

OpenStudy (anonymous):

-_- no its always TWO

OpenStudy (anonymous):

why its always two anyway =.= !

hartnn (hartnn):

really? what if base is 2 ??

OpenStudy (waleed_imtiaz):

no its 11....... :D

OpenStudy (anonymous):

o_o'

hartnn (hartnn):

11 in which base ?? i think,no.

OpenStudy (anonymous):

oh GoD >_< im outta here X[

OpenStudy (anonymous):

It depends of course on precisely what meaning you have ascribed to "+"

OpenStudy (waleed_imtiaz):

hehe.... okay its 2 always.....

OpenStudy (anonymous):

and i thought i was an easy one =_='

hartnn (hartnn):

its the question that no one here has unique answer!

OpenStudy (anonymous):

Mathematicians are experts in making mountains out of molehills....

OpenStudy (anonymous):

okie bye bye all :l i gtg i got business .. i mean homeworks to Do :B

OpenStudy (anonymous):

XD haaa byzz ^_^

hartnn (hartnn):

bye, keep on asking such impossible questions........

OpenStudy (bahrom7893):

If you define 1 to be 3 and + to be *, then the answer is 9

OpenStudy (unklerhaukus):

1 +1 = 1 +1 -1 = -1 1 = -1